Consider the following matrix: 1-5 -5 -1 3 A-1 5 5 2 -4 2-10-10 0 4 Give a set of vectors that spans each of im(A) and null(A) Number of Vectors: 1 Spanning set for im(A) Number of Vectors: 1 Spanning set for null(A)30
Answer #1

matrix is

1 -5 -5 -1 3
-1 5 5 2 -4
2 -10 -10 0 4

convert into Reduced Row Eschelon Form...

Add (1 * row1) to row2

1 -5 -5 -1 3
0 0 0 1 -1
2 -10 -10 0 4


Add (-2 * row1) to row3

1 -5 -5 -1 3
0 0 0 1 -1
0 0 0 2 -2


Add (-2 * row2) to row3

1 -5 -5 -1 3
0 0 0 1 -1
0 0 0 0 0


Add (1 * row2) to row1

1 -5 -5 0 2
0 0 0 1 -1
0 0 0 0 0

there are two pivot entry at first and fourth column

so image(A)=\large {\color{Red} \begin{pmatrix}1\\ -1\\ 2\end{pmatrix},\begin{pmatrix}-1\\ 2\\ 0\end{pmatrix}}

\large \begin{pmatrix}1&-5&-5&0&2\\ 0&0&0&1&-1\\ 0&0&0&0&0\end{pmatrix}\begin{pmatrix}a\\ b\\ c\\ d\\ e\end{pmatrix}=0

\large a-5b-5c+2e=0....................a=5b+5c-2e

\large d-e=0....................................d=e

\large b=b

\large c=c

\large e=e

\large \begin{pmatrix}a\\ \:b\\ \:c\\ \:d\\ \:e\end{pmatrix}=b\begin{pmatrix}5\\ 1\\ 0\\ 0\\ 0\end{pmatrix}+c\begin{pmatrix}5\\ 0\\ 1\\ 0\\ 0\end{pmatrix}+e\begin{pmatrix}-2\\ 0\\ 0\\ 1\\ 1\end{pmatrix}

null space(A)= \large {\color{Red} \begin{pmatrix}5\\ \:1\\ \:0\\ \:0\\ \:0\end{pmatrix}\begin{pmatrix}5\\ \:0\\ \:1\\ \:0\\ \:0\end{pmatrix}\begin{pmatrix}-2\\ \:0\\ \:0\\ \:1\\ \:1\end{pmatrix}}